OpenStudy (anonymous):

2x-3=-10,when x=4, what is the value of y?

OpenStudy (anonymous):

there is no y

OpenStudy (anonymous):

3y sorry

OpenStudy (anonymous):

substitute 4 into x. 2(4) - 3y = 10 8 - 3y = 10 -3y = 10 - 8 -3y = -2

OpenStudy (anonymous):

10 - 8 is positive 2

OpenStudy (anonymous):

sorry yeah

OpenStudy (anonymous):

How could I put that on a fraction?

OpenStudy (anonymous):

you solve for y -3y = 2

OpenStudy (anonymous):

y = -2/3

OpenStudy (anonymous):

Wait wasn't 2positive? Samyflashy

OpenStudy (anonymous):

So the whole fraction will be negative right?

OpenStudy (anonymous):

yeah but to solve for y